I have problem with 6 Martin MAC500 heads. All six have issue with tili it is not moving at all only shaking. With the small push head starts to move and stops again after few seconds - it will not go thru the reset as it it cannot reach its end position. I have checked them by conecting cable from main PCB of working unit to the small PCB on the arm of the head that I am trying to fix . As result head started tilting normaly.I have also awapped the drivers between the working and non working unit and it still did change anything. I have spent a lot of time on this head and I am stuck now. Any help will great.